# Reconstruction of frequency-localized functions from pointwise samples via least squares and deep learning

### v1

@@ 5.1 A preliminary result for the main theorems
If we accept Proposition 5.8, then the rationale behind the choice of $\kappa^{\star}=(2\,\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}$ in (3.2) can be seen from (5.20). Specifically, by ensuring $m\geq c_{\delta}\kappa^{\star}\log(\texttt{\#}\Lambda/\beta)$, the proof of Theorem 3.1 can now be completed through a combination of Theorem 5.7 and (5.2). ∎
It remains for us to establish Proposition 5.8 next.
Proof of Proposition 5.8.
We begin with an outline and a preliminary result. To derive an estimate for $\kappa(\mathcal{S}_{\Lambda})$, where $\Lambda=\Lambda^{\rm HC}_{n-1}=\{\vec{\nu}\in\mathbb{N}_{0}^{d}:\prod_{k=1}^{d}(\nu_{k}+1)\leq n\}$ and $d=1,2,3$, we utilize an induction strategy on dimension, as employed in the proof of [3, Proposition 5.13] (also [14, Lemma 3.3]). Our induction step will end at $d=3$. For the base case $d=1$, we make use of Proposition 5.1 to provide an upper bound for $\kappa(\mathcal{S}_{\Lambda})$. For the induction step, we observe from the slicing property (2.9) that each $\Lambda_{k}$ is a $(d-1)$-dimensional hyperbolic cross of order $\big{\lfloor}\frac{n}{k+1}\big{\rfloor}$, i.e.
@@
is a $(d-1)$-dimensional hyperbolic cross of exactly order $n$. From there, we rely on the auxiliary yet crucial result stated below, which compares the sizes of horizontal slices in two- and three-dimensional hyperbolic cross index sets. Its proof is given at the conclusion of this subsection.
@@
Returning to the proof of Proposition 5.8, we note that $\gamma(\mathsf{w})>1$ for $\mathsf{w}\geq 1$. Moreover
@@
$\kappa(\mathcal{S}_{\Lambda})=\|\varphi_{\mathsf{w},0}\|^{2}_{L^{\infty}([-1,1])}\leq\frac{4\mathsf{w}}{\pi}\leq 2^{\gamma(\mathsf{w})}=(\sqrt{2}\,\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})},$ (5.25)
@@
$\kappa(\mathcal{S}_{\Lambda})=\sum_{k=0}^{n-1}\|\varphi_{\mathsf{w},k}\|^{2}_{L^{\infty}([-1,1])}\leq\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(k+1)^{\gamma(\mathsf{w})}\leq\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(2k+1)^{\gamma(\mathsf{w})}.$ (5.26)
@@
$\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(2k+1)^{\gamma(\mathsf{w})}\leq\Big{(}\sum_{k=0}^{n-1}2k+1\Big{)}^{\gamma(\mathsf{w})}=(n^{2}-1)^{\gamma(\mathsf{w})}\leq(2n^{2})^{\gamma(\mathsf{w})}=(\sqrt{2}\,\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}.$ (5.27)
It is enough to validate the first inequality in (5.27). Then on account of Pascal’s identity [37, Chapter 6.4, Theorem 2] and (5.24), we have
@@
$\geq\sum_{l=0}^{\gamma(\mathsf{w})-1}{\gamma(\mathsf{w})\choose l}\Big{(}\sum_{k=1}^{n-1}2k+1\Big{)}^{l}+\sum_{k=1}^{n-1}(2k+1)^{\gamma(\mathsf{w})}$
@@
Turning to the case $d=2$ or $d=3$, we have that $\Lambda=\Lambda^{\rm HC}_{n-1}$ is a two-dimensional or three-dimensional hyperbolic cross, where each $\Lambda_{k}$ is a one-dimensional or two-dimensional hyperbolic cross, respectively. In what follows, we will employ an argument applicable for both the case $d=2$ when $n\geq 2$ and for $d=3$ when $n\geq 26$. For the latter case, we additionally assume that (5.20) already holds for $d=2$ for all $n\in\mathbb{N}$. The separate case of $d=2$ and $n=1$ will be addressed at the end.
Let $d=2$ and $n\geq 2$. By applying (5.20) to each one-dimensional hyperbolic cross $\Lambda_{k}$ and invoking Proposition 5.1, we obtain
$\kappa(\mathcal{S}_{\Lambda})$ $=\sum_{k=0}^{n-1}\|\varphi_{\mathsf{w},k}\|^{2}_{L^{\infty}([-1,1])}\sum_{j\in\Lambda_{k}}\|\varphi_{\mathsf{w},j}\|^{2}_{L^{\infty}([-1,1])}$
@@
Similarly, let $d=3$ and $n\geq 26$. Assuming that (5.20) holds for all two-dimensional hyperbolic cross index sets $\Lambda_{k}$, we can write
As with the case $d=1$, we aim to dominate the right-hand sides of (5.2), (5.2) using the quantity $(\sum_{k=0}^{n-1}(2k+1)\,(2\,\texttt{\#}\Lambda_{k})^{2})^{\gamma(\mathsf{w})}$. For brevity, we denote $A_{k}:=2\,\texttt{\#}\Lambda_{k}$. Then
@@
where the last inequality is due to (5.24). Putting back $A_{k}=2\,\texttt{\#}\Lambda_{k}$, and combining (5.30), (5.31) together with (5.2) or (5.2) yields
@@
which is (5.20) for $d=2$ when $n\geq 2$ and for $d=3$ when $n\geq 26$.
It remains to check the validity of (5.20) when $d=2$ and $n=1$. In this case, $\Lambda=\{(0,0)\}$. Then performing a similar calculation as in (5.25) yields
$\kappa(\mathcal{S}_{\Lambda})=\|\varphi_{\mathsf{w},(0,0)}\|^{2}_{L^{\infty}([-1,1]^{2})}=\|\varphi_{\mathsf{w},0}\|^{4}_{L^{\infty}([-1,1])}\leq\frac{16\mathsf{w}^{2}}{\pi^{2}}\leq 2^{2\gamma(\mathsf{w})}=(2\,\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}.$
With this, we close the induction loop and conclude (5.20) for $d=1,2$ for all $n\in\mathbb{N}$ and for $d=3$ for all $n\geq 26$. ∎
Proof of Proposition 5.9.
We rely on the following cardinality estimate of the hyperbolic cross.
Lemma 5.10 (See [13, Theorem 3.5] and [3, Lemma B.3]).
#$\texttt{\#}\Lambda^{\rm HC}_{n-1}<\frac{n(\log n+d\log 2)^{d-1}}{(d-1)!}.$
An estimate on $n_{\star}(d)$ can be derived, for example, from the calculations presented in the proof of [13, Theorem 2.4]. In particular, we can take $n_{\star}(2)=1$. We proceed to prove Proposition 5.9 for $d=2$. In this case, $\Lambda=\Lambda^{\rm HC}_{n-1}$ is a two-dimensional hyperbolic cross where each slice $\Lambda_{k}$, for $k=0,\dots,n-1$, is a one-dimensional hyperbolic cross. It also follows from (5.22) that #$\texttt{\#}\Lambda_{0}=n$. Therefore, proving the proposition amounts to proving
$\sum_{k=1}^{n-1}(2k+1)\,(\texttt{\#}\Lambda_{k})^{2}\geq n.$
However, it can be seen that the right-hand side term above is at least
$\sum_{k=1}^{n-1}(2k+1)\,(\texttt{\#}\Lambda_{k})^{2}\geq\sum_{k=1}^{n-1}(2k+1)=n(n-1)+n-1=n^{2}-1\geq n,$
when $n\geq 2$, and we are done with $d=2$.
Turning to $d=3$, we obtain from the Cauchy-Schwarz inequality that
$\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})^{2}(2k+1)\geq\frac{1}{n}\big{(}\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})\sqrt{2k+1}\big{)}^{2}.$ (5.33)
Note that in this case, each $\Lambda_{k}$ is a two-dimensional hyperbolic cross. Further, we claim:
Claim 5.11.
$\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})\sqrt{2k+1}\geq\texttt{\#}\Lambda_{0}$, whenever $n\geq 26$.
We demonstrate this. First, since #$\texttt{\#}\Lambda_{k}\geq 1$ for all $k\leq n-1$, we have
$\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})\sqrt{2k+1}\geq\sum_{k=1}^{n-1}\sqrt{2k+1}\geq\int_{0}^{n-1}(2t+1)^{1/2}\,{\rm d}t\geq\frac{(2n-1)^{3/2}-1}{3}.$ (5.34)
Second, from Lemma 5.10 and the fact that $n_{\star}(2)=1$, we have #$\texttt{\#}\Lambda_{0}\leq n(\log n+2\log 2)$, whenever $n\geq 1$. Third,
$\frac{(2n-1)^{3/2}-1}{3}\geq n(\log n+2\log 2),$ (5.35)
as soon as $n\geq 26$. It is now apparent that Claim 5.11 follows from (5.34), (5.35).
Combining (5.33) and Claim 5.11, we deduce for $d=3$ and $n\geq 26$,
$\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})^{2}(2k+1)\geq\frac{1}{n}\big{(}\sum_{k=1}^{n-1}(\texttt{\#}\Lambda_{k})\sqrt{2k+1}\big{)}^{2}\geq\frac{(\texttt{\#}\Lambda_{0})^{2}}{n},$
which is the wanted conclusion. ∎
@@
The culmination of Step 1 and Step 2 is Proposition 5.12, the main preliminary result of this subsection, which will be presented shortly. The outcome of Step 3 will be our second main theorem, Theorem 3.2, with its proof provided in Subsection 5.4. The initial steps rely on the fact that Legendre polynomials can effectively approximate Slepian basis functions. Thus, we begin with an overview of Legendre polynomials.
@@
such that $\max_{x\in[-1,1]}|\widetilde{P}_{k}(x)|=1$. The family $\{\widetilde{P}_{k}\}_{k\in\mathbb{N}_{0}}$ forms an orthogonal basis in $L^{2}_{\bf u}([-1,1])$. Moreover, as shown in, e.g., [3, Subsection 2.2.2],
@@
It is known that the normalized Legendre polynomials $\{P_{k}\}_{k\in\mathbb{N}_{0}}$ in one dimension can be effectively approximated by ReLU neural networks. A concrete example is provided in [30, Proposition 2.11] (see also Proposition A.7). In the case of $d$ dimensions, we frequently rely on the tensorization definition (5.37) and the fundamental network calculus operations outlined in Appendix A, such as concatenation, parallelization, and product approximation, to develop approximating NNs for normalized Legendre polynomials. Although these constructions are well-established in the literature, we include them for completeness and convenient reference as the reader progresses through the material.
@@
Next, we utilize the network construction ×~$\tilde{\times}_{\varepsilon,B}$ in Proposition A.1, with $B=2n^{\gamma(\mathsf{w})}+1$ (obtained from (5.48)), as well as the concatenation and parallelization operations, denoted as $\odot$ and ${\rm P}$, from Definitions A.2 and A.3 respectively, to construct
@@
It follows the construction that
@@
see also [31, Equation (2.47)]. Since $L^{\infty}([-1,1])$ is a Banach space, the absolute convergence of the series $\sum_{j=0}^{\infty}\beta_{j}^{k}P_{j}$, i.e.
@@
will guarantee the existence of the limit $\sum_{j=0}^{\infty}\beta_{j}^{k}P_{j}$ as a function in $L^{\infty}([-1,1])$ [20, Theorem 5.1]. Once held, (5.57) together with (5.56), and the continuity of $\sum_{j=0}^{N}\beta_{j}^{k}P_{j}$ and $\varphi_{\mathsf{w},k}$, yields
@@
on $[-1,1]$. We proceed to demonstrate (5.57) for all $k\in\Lambda$. By [31, Theorem 7.2], if $j\geq 2(\lfloor e\mathsf{w}\rfloor+1)$, then
@@
Let $j=0,\dots,N$. A result derived from [30, Proposition 2.11], and stated as Proposition A.7 in Appendix A, guarantees for $j=1,\dots,N$, the existence of an NN $\Phi^{j}_{\varepsilon}$ with both input dimension and output dimension $d^{j}_{L_{j}}$ equal to $1$, such that
@@
and for $j=0$, an NN $\Phi^{0}_{\varepsilon}$ satisfying ${\rm R}(\Phi^{0}_{\varepsilon})=P_{0}=1$ on $[-1,1]$. Moreover,
@@
To simplify the presentation, we adopt the convention introduced at the beginning of Subsection 5.2. Namely, we use $\Lambda$ to denote $\Lambda^{\rm HC}_{n-1}$, and let $K=\texttt{\#}\Lambda=\texttt{\#}\Lambda^{\rm HC}_{n-1}$. For $\vec{\nu}\in\Lambda$, we write $\varphi_{\mathsf{w},\vec{\nu}}=\varphi_{\mathsf{w},\Pi(k)}$, where $\Pi:\{1,\dots,K\}\to\Lambda$ is a bijection. Now let $g\in\mathcal{S}_{\Lambda}$, i.e.
$g=\sum_{k=1}^{K}b_{k}\varphi_{\mathsf{w},\Pi(k)},$ (5.68)
for $b_{k}:=\frac{1}{2}\int_{-1}^{1}g(x)\varphi_{\mathsf{w},\Pi(k)}(x)\,{\rm d}x$. According to Proposition 5.12, for a given $\varepsilon>0$ and each Slepian basis function $\varphi_{\mathsf{w},\Pi(k)}$, there exists an NN $\Psi_{\varepsilon}^{k}$ such that [^4] Given that $N_{\star}$, in (5.38), depends on $\varepsilon$, we will omit its reference in the subscript and simply write $\Psi^{k}_{\varepsilon}$ going forward.
$\|\varphi_{\mathsf{w},k}-{\rm R}(\Psi_{\varepsilon}^{k})\|_{L^{\infty}([-1,1]^{d})}\leq B(d,n)\varepsilon.$ (5.69)
@@
$\|g-{\rm R}(\Psi_{g;\varepsilon})\|_{L^{\infty}([-1,1]^{d})}$ $\leq\sum_{k=1}^{K}|b_{k}|\|\varphi_{\mathsf{w},k}-{\rm R}(\Psi_{\varepsilon}^{k})\|_{L^{\infty}_{\bf u}([-1,1]^{d})}$
$\leq\sqrt{K}B(d,n)\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}\varepsilon,$ (5.71)
from (5.69) and the fact that $\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}=(\sum_{k=1}^{K}|b_{k}|^{2})^{1/2}$. It also follows from Proposition 5.12 and another application of Proposition A.6 that
$\begin{split}L(\Psi_{g;\varepsilon})&\leq C\big{(}(1+\log_{2}N_{\star})(N_{\star}+\log_{2}(N_{\star}/\varepsilon))+(1+\log_{2}(M(d,n)/\varepsilon)\big{)},\\ W(\Psi_{g;\varepsilon})&\leq CK\big{(}(N_{\star}^{2}+N_{\star})(N_{\star}+\log_{2}(N_{\star}/\varepsilon))+(1+\log_{2}(M(d,n)/\varepsilon)\big{)}.\end{split}$ (5.72)
@@
Recall the matrix $A\in\mathbb{R}^{m\times K}$, introduced in (5.16), where $A_{jk}=\dfrac{1}{\sqrt{m}}\varphi_{\mathsf{w},\Pi(k)}(\vec{y}_{j})$. Since the Slepian basis is an orthonormal set, the columns of $A$ are linearly independent. Consider $\tilde{A}\in\mathbb{R}^{m\times K}$, such that $\tilde{A}_{jk}=\dfrac{1}{\sqrt{m}}{\rm R}(\Psi_{\varepsilon}^{k})(\vec{y}_{j})$. Then it follows from (5.69) that the columns of $\tilde{A}$ are also linearly independent for sufficiently small $\varepsilon$. Indeed, let $\varepsilon\in(0,1/(\sqrt{K}B(d,n)))$. If $0=\sum_{k=1}^{K}b_{k}{\rm R}(\Psi^{k}_{\varepsilon})$ for some not-all-zero scalars $\{b_{k}\}_{k=1}^{K}$, then
$\sum_{k=1}^{K}b_{k}(\varphi^{k}_{\varepsilon}-{\rm R}(\Psi^{k}_{\varepsilon}))=\sum_{k=1}^{K}b_{k}\varphi^{k}_{\varepsilon},$

### v2

@@ 5.1 A preliminary result for the main theorems
The proof of Proposition 5.8 utilizes an induction argument on dimension, a strategy employed in the proof of [adcock2022sparse, Proposition 5.13] (also [chkifa2015discrete, Lemma 3.3]). For the base case $d=1$, we make use of Proposition 5.1 to provide an upper bound for $\kappa(\mathcal{S}_{\Lambda})$. For the induction step, we observe from the slicing property (2.9) that each $\Lambda_{k}$ is a (d1)(d-1)-dimensional hyperbolic cross of order $\big\lfloor\frac{n}{k+1}\big\rfloor$, i.e.
@@
is a (d1)(d-1)-dimensional hyperbolic cross of exactly order $n$. From there, we rely on the auxiliary yet crucial result stated below, which compares the sizes of horizontal slices in two- and three-dimensional hyperbolic cross index sets.
@@
Proof.
We first prove the proposition for $d=2$. In this case, $\Lambda=\Lambda^{\rm HC}_{n-1}$ is a two-dimensional hyperbolic cross where each slice $\Lambda_{k}$, for $k=0,\dots,n-1$, is a one-dimensional hyperbolic cross. From (5.18), we have $\texttt{\#}\Lambda_{0}=n$. Thus, proving the proposition amounts to proving
$\sum_{k=1}^{n-1}(2k+1)\,(\texttt{\#}\Lambda_{k})^{2}\geq n.$
However, this is immediate, since
$\sum_{k=1}^{n-1}(2k+1)\,(\texttt{\#}\Lambda_{k})^{2}\geq\sum_{k=1}^{n-1}(2k+1)=n^{2}-1\geq n,$
when $n\geq 2$, and we are done with $d=2$. Turning to $d=3$, we observe that
$\sum_{k=1}^{n-1}(2k+1)(\texttt{\#}\Lambda_{k})^{2}\geq 3(\texttt{\#}\Lambda_{1})^{2}\geq 3\Big\lfloor\frac{n}{2}\Big\rfloor^{2}.$ (5.20)
Note that in this case, each $\Lambda_{k}$ is a two-dimensional hyperbolic cross. Here, the last inequality in (5.20) holds since $\Lambda_{1}$ contains the set
$\Big\{(0,j):0\leq j\leq\Big\lfloor\frac{n}{2}\Big\rfloor-1\Big\}.$
At the same time, we can write
$\Lambda_{0}=\bigcup_{k=0}^{n-1}\Big\{(k,j),0\leq j\leq\frac{n}{k+1}-1\Big\},$
which gives
$\texttt{\#}\Lambda_{0}\leq\sum_{k=0}^{n-1}\frac{n}{k+1}\leq n\Big(1+\int_{1}^{n}\frac{1}{x}\,{\rm d}x\Big)=n(1+\log(n)).$ (5.21)
Therefore, if
$3\Big\lfloor\frac{n}{2}\Big\rfloor^{2}\geq n(1+\log(n))^{2},$ (5.22)
then the desired conclusion (5.19) follows from (5.20) and (5.21). A direct check shows that (5.22) holds for all $n\geq 26$, which completes the proof. ∎
We now proceed with the proof of Proposition 5.8.
Proof of Proposition 5.8.
To begin, we note that $\gamma(\mathsf{w})>1$ for $\mathsf{w}\geq 1$. Moreover
@@
$\kappa(\mathcal{S}_{\Lambda})=\|\varphi_{\mathsf{w},0}\|^{2}_{L^{\infty}([-1,1])}\leq\frac{4\mathsf{w}}{\pi}\leq 2^{\gamma(\mathsf{w})}\leq(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})},$ (5.24)
@@
$\kappa(\mathcal{S}_{\Lambda})\leq\sum_{k=0}^{n-1}\|\varphi_{\mathsf{w},k}\|^{2}_{L^{\infty}([-1,1])}\leq\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(k+1)^{\gamma(\mathsf{w})}\leq\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(2k+1)^{\gamma(\mathsf{w})}.$ (5.25)
@@
$\frac{4\mathsf{w}}{\pi}+\sum_{k=1}^{n-1}(2k+1)^{\gamma(\mathsf{w})}\leq\Big(\sum_{k=0}^{n-1}(2k+1)\Big)^{\gamma(\mathsf{w})}=n^{2\gamma(\mathsf{w})}\leq(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}.$ (5.26)
It is enough to validate the first inequality in (5.26). Then on account of Pascal’s identity [rosen1999discrete, Chapter 6.4, Theorem 2] and (5.23), we have
@@
When $d=2$ and $n=1$, we have $\Lambda=\{(0,0)\}$. Then a similar calculation as in (5.24) yields
$\kappa(\mathcal{S}_{\Lambda})=\|\varphi_{\mathsf{w},(0,0)}\|^{2}_{L^{\infty}([-1,1]^{2})}=\|\varphi_{\mathsf{w},0}\|^{4}_{L^{\infty}([-1,1])}\leq\frac{16\mathsf{w}^{2}}{\pi^{2}}\leq 2^{2\gamma(\mathsf{w})}=(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})},$
which establishes (5.16). When $d=2$ and $n\geq 2$, we apply (5.16) to each one-dimensional hyperbolic cross $\Lambda_{k}$ (5.17), (5.18) and invoke Proposition 5.1, obtaining
$\kappa(\mathcal{S}_{\Lambda})$ $\leq\sum_{k=0}^{n-1}\|\varphi_{\mathsf{w},k}\|^{2}_{L^{\infty}([-1,1])}\sum_{j\in\Lambda_{k}}\|\varphi_{\mathsf{w},j}\|^{2}_{L^{\infty}([-1,1])}$
@@
As with the case $d=1$, we aim to bound the right-hand side of (5.2) by $(\sum_{k=0}^{n-1}(2k+1)\,(2\texttt{\#}\Lambda_{k})^{2})^{\gamma(\mathsf{w})}$. For brevity, we denote $A_{k}:=2\texttt{\#}\Lambda_{k}$. Then
@@
where the last inequality is due to (5.23). Putting back $A_{k}=2\texttt{\#}\Lambda_{k}$, and combining (5.28), (5.29) together with (5.2) yields
@@
which is (5.16) when $d=2$ and $n\geq 2$.
Finally, when $d=3$ and $n\geq 26$, we adopt arguments similar to those used for $d=2$ and $n\geq 2$. Namely, we first apply (5.16) to each two-dimensional hyperbolic cross $\Lambda_{k}$ (5.17), (5.18), and invoke Proposition 5.1, to obtain
$\leq\frac{4\mathsf{w}}{\pi}(2\texttt{\#}\Lambda_{0})^{2\gamma(\mathsf{w})}+\sum_{k=1}^{n-1}(k+1)^{\gamma(\mathsf{w})}(2\texttt{\#}\Lambda_{k})^{2\gamma(\mathsf{w})},$ (5.31)
which matches the form of (5.2). From this point on, the argument proceeds exactly as in the case $d=2$ and $n\geq 2$, with the sole modification being the application of Proposition 5.9 for $d=3$ and $n\geq 26$. We therefore omit further details. ∎
5.3 Proof of Theorem 3.1
Building on Proposition 5.8, if we take
$m\geq((1-\delta)\log(1-\delta)+\delta)^{-1}(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\log(\texttt{\#}\Lambda/\beta),$
then we guarantee (5.15) from Theorem 5.7. From there, the theory outlined in [adcock2022sparse, Section 5], particularly [adcock2022sparse, Corollary 5.9], demonstrates that every $f\in\mathcal{C}([-1,1]^{d})$ has a unique least squares approximation $f^{\natural}$ in (2.8) that satisfies
$<\Big(1+\frac{1}{\sqrt{1-\delta}}\Big)\inf_{g\in\mathcal{S}_{\Lambda}}\|f-g\|_{L^{\infty}([-1,1]^{d})}+\frac{1}{\sqrt{1-\delta}}\|\vec{e}\|_{2},$ (5.32)
thereby completing the proof of Theorem 3.1. ∎
@@
The culmination of Step 1 and Step 2 is Proposition 5.10, the main preliminary result of this subsection, which will be presented shortly. The outcome of Step 3 will be our second main theorem, Theorem 3.3, with its proof provided in Section 5.5. The initial steps rely on the fact that Legendre polynomials can effectively approximate Slepian basis functions. Thus, we begin with an overview of Legendre polynomials.
@@
such that $\max_{x\in[-1,1]}|\widetilde{P}_{k}(x)|=1$. The family $\{\widetilde{P}_{k}\}_{k\in\mathbb{N}_{0}}$ forms an orthogonal basis in $L^{2}_{\bf u}([-1,1])$. Moreover, as shown in, e.g., [adcock2022sparse, Section 2.2.2],
@@
It is known that the normalized Legendre polynomials $\{P_{k}\}_{k\in\mathbb{N}_{0}}$ in one dimension can be effectively approximated by ReLU neural networks. A concrete example is provided in [opschoor2022exponential, Proposition 2.11] (see also Proposition A.7). In the case of $d$ dimensions, we frequently rely on the tensorization definition (5.34) and the fundamental network calculus operations outlined in Appendix A, such as concatenation, parallelization, and product approximation, to develop approximating NNs for normalized Legendre polynomials. Although these constructions are well-established in the literature, we include them for completeness and convenient reference as the reader progresses through the material.
@@
Next, we utilize the network construction $\tilde{\times}_{\varepsilon,B}$ in Proposition A.1, with $B=2n^{\gamma(\mathsf{w})}+1$ (obtained from (5.45)), as well as the concatenation and parallelization operations, denoted as $\odot$ and $\rm P$}, from Definitions A.2 and A.3 respectively, to construct
@@
It follows from the construction that
@@
see also [osipov2013prolate, Equation (2.47)]. Since $L^{\infty}([-1,1])$ is a Banach space, the absolute convergence of the series $\sum_{j=0}^{\infty}\beta_{j}^{k}P_{j}$, i.e.
@@
will guarantee the existence of the limit $\sum_{j=0}^{\infty}\beta_{j}^{k}P_{j}$ as a function in $L^{\infty}([-1,1])$ [folland1999real, Theorem 5.1]. Once held, (5.54) together with (5.53), and the continuity of $\sum_{j=0}^{N}\beta_{j}^{k}P_{j}$ and $\varphi_{\mathsf{w},k}$, yields
@@
on [1,1][-1,1]. We proceed to demonstrate (5.54) for all $k\in\Lambda$. By [osipov2013prolate, Theorem 7.2], if $j\geq 2(\lfloor e\mathsf{w}\rfloor+1)$, then
@@
Let $j=0,\dots,N$. A result derived from [opschoor2022exponential, Proposition 2.11], and stated as Proposition A.7 in Appendix A, guarantees for $j=1,\dots,N$, the existence of an NN $\Phi^{j}_{\varepsilon}$ with both input dimension and output dimension $d^{j}_{L_{j}}$ equal to $1$, such that
@@
and for $j=0$, an NN $\Phi^{0}_{\varepsilon}$ satisfying ${\rm R}(\Phi^{0}_{\varepsilon})=P_{0}=1$ on [1,1][-1,1]. Moreover,
@@
Let $g\in\mathcal{S}_{\Lambda}$, i.e.
$g=\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}\varphi_{\mathsf{w},\vec{\nu}},$ (5.65)
for $b_{\vec{\nu}}:=\frac{1}{2}\int_{-1}^{1}g(x)\varphi_{\mathsf{w},\vec{\nu}}(x)\,{\rm d}x$. According to Proposition 5.10, for a given $\varepsilon>0$ and each Slepian basis function $\varphi_{\mathsf{w},\vec{\nu}}$, there exists an NN $\Psi_{\varepsilon}^{\vec{\nu}}$ such that [^5] Given that $N_{\star}$, in (5.35), depends on $\varepsilon$, we will omit its reference in the subscript and simply write $\Psi^{\vec{\nu}}_{\varepsilon}$ going forward.
$\|\varphi_{\mathsf{w},\vec{\nu}}-{\rm R}(\Psi_{\varepsilon}^{\vec{\nu}})\|_{L^{\infty}([-1,1]^{d})}\leq B(d,n)\varepsilon.$ (5.66)
@@
$\|g-{\rm R}(\Psi_{g;\varepsilon})\|_{L^{\infty}([-1,1]^{d})}$ $\leq\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|\|\varphi_{\mathsf{w},\vec{\nu}}-{\rm R}(\Psi_{\varepsilon}^{\vec{\nu}})\|_{L^{\infty}_{\bf u}([-1,1]^{d})}$
$\leq\sqrt{\texttt{\#}\Lambda}\,B(d,n)\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}\varepsilon,$ (5.68)
from (5.66) and the fact that $\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}=(\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|^{2})^{1/2}$. It also follows from Proposition 5.10 and another application of Proposition A.6 that
$\begin{split}L(\Psi_{g;\varepsilon})&\leq C\big((1+\log_{2}N_{\star})(N_{\star}+\log_{2}(N_{\star}/\varepsilon))+(1+\log_{2}(M(d,n)/\varepsilon)\big),\\ W(\Psi_{g;\varepsilon})&\leq C\texttt{\#}\Lambda\big((N_{\star}^{2}+N_{\star})(N_{\star}+\log_{2}(N_{\star}/\varepsilon))+(1+\log_{2}(M(d,n)/\varepsilon)\big).\end{split}$ (5.69)
@@
Recall the matrix $A\in\mathbb{R}^{m\times\texttt{\#}\Lambda}$, introduced in (5.13), where $A_{j,\vec{\nu}}=\dfrac{1}{\sqrt{m}}\varphi_{\mathsf{w},\vec{\nu}}(\vec{y}_{j})$. Since the Slepian basis is an orthonormal set, the columns of $A$ are linearly independent. Consider $\tilde{A}\in\mathbb{R}^{m\times\texttt{\#}\Lambda}$, such that $\tilde{A}_{j,\vec{\nu}}=\dfrac{1}{\sqrt{m}}{\rm R}(\Psi_{\varepsilon}^{\vec{\nu}})(\vec{y}_{j})$. Then it follows from (5.66) that the columns of $\tilde{A}$ are also linearly independent for sufficiently small $\varepsilon$. Indeed, let $\varepsilon\in(0,1/(\sqrt{\texttt{\#}\Lambda}\,B(d,n)))$. If $0=\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon})$ for some not-all-zero scalars $\{b_{\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$, then
$\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}(\varphi^{\vec{\nu}}_{\varepsilon}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon}))=\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}\varphi^{\vec{\nu}}_{\varepsilon},$